OpenStudy (anonymous):

Solve each equation. 9(z-3) = 12z 7y+5 = 6y + 11 5w+8 - 12w = 16 - 15w 3(x+1) = 2(x+11)

OpenStudy (anonymous):

Are these a system of equations or are just solve each equation for the included variable?

OpenStudy (anonymous):

Each one is a different equation.

OpenStudy (***[isuru]***):

Hi, here's the basic to solve them.... u first have to remove the brackets of each equation by multiplying them by the given number.. for an example if there was 9(x + 3) u have to simplify it to (9x) + (9x3) = 9x+27 then u can add or subtract like terms in the equation... which mean for an instance... 4x + 5 + x +6 here u can add 4x to x and get 5x .... but u cant add 4x to 5,6... then u can get the expression 5x + 11 and finally to solve them.. u have keep doing this until u have only 2 terms in the equation... one is known and other one is unknown... 5y = 15 like that then u can divide both sides by a suitable no and get the value for the unknown term.... 5y/5 = 15/5 y = 3 so to help ya... wt will be ur 1st equation if u remove the brackets of it as I mentioned above ?